OpenStudy (anonymous):

PLEASE HELP! I GIVE MEDALS! Samantha loves to download new music. She originally had 52 songs but plans to purchase 2 new songs each week. She wants to know how many songs she will have after 93 weeks. Which equation should she use? a = 93 + 2(52 - 1) a = 93 + 2(52 + 1) a = 52 + 2(93 + 1) a = 52 + 2(93 - 1)

OpenStudy (anonymous):

93+2(52-1) I think hope it helps!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

A new video game is expected to sell 120 copies the first hour at a local game store. After that, the sales will follow the function s(x) = 15(x - 1) where x is the number of hours. What is the function that shows total sales, including the first hour? s(x) = 15(x - 1) - 120 s(x) = 120x - 15 s(x) = 120 + 15(x - 1) s(x) = 120 - 15(x - 1)

OpenStudy (preetha):

Can you look at the above response and try to figure it out? :)

OpenStudy (anonymous):

Do you mean it is c?